OpenStudy (anonymous):

Super-uper urgent basic "differentiate with respect to x" problem: sqrt of [3 - 4x] Since my booklet says formula is x^n = n*x^(n-1) I got 0.5*(3-4x)^(-1/2) But my teacher wrote in my failed exam that correct answer is: 0.5*(3-4x)^(-1/2) , (-4) I don't get it at all why is this so

OpenStudy (anonymous):

You did that part right, but you need to apply the chain rule as well. You want \[f^{\prime}(x)= \frac{df}{du}\frac{du}{dx}\]

OpenStudy (anonymous):

\[u=3-4x\] so \[f(u)=(u)^{1/2}\] and \[f^{\prime}(u) = \frac{1}{2}u^{-1/2}\] But you are instructed to find f(x).

OpenStudy (anonymous):

\[\frac{du}{dx}=-4\] so \[f^{\prime} = \frac{1}{2}u^{-1/2}(-4)\] which becomes \[f^{\prime}(x) =-2(3-4x)^{-1/2}\] when you back substitute what u was

OpenStudy (anonymous):

okay so after ive done the process of finding derivative i need to use chain's rule after that. im still trying to figure out how it works with your explanation hang on

OpenStudy (anonymous):

du/dx is the coefficient of the variable we are differentiating with respect to, in this case x?

OpenStudy (anonymous):

what is df/du

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

with the substitution \[f(u)=u^{1/2}\] so df/du is \[\frac{1}{2}u^{-1/2}\]

OpenStudy (anonymous):

df/du is the thing i found and thought was the final answer? so i get df/du by differentiating the specified thing with a variable and multiply this by the coefficient of the variable we are differentiating with respect to?

OpenStudy (anonymous):

specified thing with variable is sqrt of 3-4x

OpenStudy (anonymous):

not the coefficient the derivative of u with respect to the original variable you are differentiating with

OpenStudy (anonymous):

So if F(x) = (stuff)^5 then the derivative would be 5(stuff)^(5-1) times the derivative of stuff with respect to x

OpenStudy (anonymous):

okay the derivative is gotten by using the specified formula for a type of case. in addition to this (df/du) i will multiply this by du/dx - which is just coefficient of x? correct me if/how im wrong with the last one if i am

OpenStudy (anonymous):

Example \[g(x) = (4x)^3\] \[g^{\prime}(x) = 3(4x)^2\cdot \frac{d}{dx}(4x)\]

OpenStudy (anonymous):

d/dx (4x) you just wrote there becomes just 4?

OpenStudy (anonymous):

Yeah in this case. because the derivative of ax is a. If instead you had something weird like \[1-x^7\] in the paretheses you would have multiplied by \[-7x^6\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Another example \[j(x) = (2x^2 +x)^4\] then the derivative would be \[j^{\prime}(x) = 4(2x^2+x)^3\cdot (4x+1)\] since \[\frac{d}{dx}(2x^2+x)=(4x+1)\]

OpenStudy (anonymous):

ah so in a bit more difficult case you have two variables or more inside the brackets. you just differentiate what is inside the brackets by the normal rules so that you get 2x^2 -> 4x and x -> 1 and this is what is now what was 4 in my exercise!

OpenStudy (anonymous):

any time when differentiating with respect to [variable] i have to do both take appropriate derivatives and then multiply them together for the final answer - except when i have e^x ?

OpenStudy (anonymous):

\[f(x) = e^{2x+5}\] u=2x+5 so \[f^{\prime}(x) = e^{2x+5}(\frac{d}{dx}(2x+5))\]

OpenStudy (anonymous):

So \[f^{\prime}=2e^{2x+5}\] The chain rule is used for when you substitute something more complicated for x.

OpenStudy (anonymous):

more complicated for x - so you mean in this case the more complicated for x is 2x+5 ?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

so in what you did i have to do something extra hmm let's see - so i take the exponent of the e^x and take the derivate of that. that results in 2+0. the 2 is multiplied by the derivative of e^x as a whole, which is just e^x (so e^2x+5 is e^2x+5)

OpenStudy (anonymous):

\[=1/2(3-4x)^-1/2(-4)\]

OpenStudy (anonymous):

Yes, exactly. This way they teach you something simpler like the derivative wrt x and then you learn how to do more complicated stuff.

OpenStudy (anonymous):

nx^(n-1)d/dx(base)

OpenStudy (anonymous):

as a fist rule i always first use formula to get dx/dy and if there is a composite (i think this is what it is, for example exponent with more than just x terms) i differentiate this composite to be multiplied with the derivative of dx/dy - if i understood right

OpenStudy (anonymous):

yes, but try to use the substitution of variables. You need to get used to substituting variables left and right and not getting confused. It takes a while.

OpenStudy (anonymous):

so the derivative of composite takes the role of du/dx

OpenStudy (anonymous):

what do you mean exactly by substitution of variables?

OpenStudy (anonymous):

yes. or just some other variable other than x if they doing df/dx

OpenStudy (anonymous):

see my reply you wiil get ypur problem

OpenStudy (anonymous):

thanks 0322 i got it but freewilly had already helped the same before :)

OpenStudy (anonymous):

Substitution of variables takes the difficult problem and writes it into something you recognize. You might not see \[ e^{5x^2 \cos(x)}\] and know how to deal with it, but you know how to deal with \[e^u\] because \[\frac{d}{du} e^u = e^u\]

OpenStudy (anonymous):

That way when you are given a monster of a problem instead of giving up you can break it into smaller doable pieces.

OpenStudy (anonymous):

yes sounds about right. in that example i must identify that the derivative dx/dy of the first of the two multiples (e^stuff being the first and cos(x) being the second) is just the expression e^(5x^2) itself. therefore i multiply that by what comes out of derivating (5x^2), that is, 10x

OpenStudy (anonymous):

since that is 10x and we have the e^(5x^2) half of the problem is solved, so 10x*1^(5x^2)

OpenStudy (anonymous):

then as we have broken this into two there's just cos(x) on the right side that becomes -sinx so the f inal answer is -sinx multiplied by the 10x* thing i wrote

OpenStudy (anonymous):

i usually do well when people explain this stuff to me with examples but in the exams that i have i tend to panic because of the variations of problems that there are :( i would have more questions if you had time about my exam area of mainly basic calculus

OpenStudy (anonymous):

thanks so much so far because this is as good or better than my good math teacher's explanations of the derivative and differentiation and chain's rule as a whole

OpenStudy (anonymous):

The derivative of \[5x^2\cos(x)\] wasn't something I wanted you to do. It was more of something that wasn't trivial like 4x. you need the product rule for this one \[e^{5x^2\cos{x}}(10x\cos(x)-5x^2\sin(x))\]

OpenStudy (anonymous):

ooh i almost forgot about this

OpenStudy (anonymous):

The best way is to understand the different rules with simple memorable examples that you can then see the rules when you are freezing up and apply them to the harder problem. I have to go, but good luck!

OpenStudy (anonymous):

is e^(5x^2)*cosx really what u wanted to type there?

OpenStudy (anonymous):

if u answer that then id be more than happy to let you go : D